# Top-Theta Attention: Sparsifying Transformers by Compensated Thresholding

### v1

@@ Appendix E Evaluation statistics
(a) LLaMA3-8B-Instruct fraction of attention elements needed
(b) LLaMA3-70B-Instruct fraction of attention elements needed
Figure 16: Generative Task Human-eval- during prefill - Tradeoff between model accuracy averaged across test samples (y-axis), and the portion of required attention elements per head (x-axis). The Top-{greek}j variants employ threshold calibrated on ARC-C dataset. All post-softmax Top-k and Top-{greek}j employ VMC, all pre-softmax variants employ both VMC and exact SDC
(a) LLaMA3-8B-Instruct fraction of V-rows needed
(b) LLaMA3-70B-Instruct fraction of V-rows needed
Figure 17: Generative Task Human-eval- during generative decoding - Tradeoff between model accuracy averaged across test samples (y-axis), and the portion of required V-rows per group of heads (x-axis). The Top-{greek}j variants employ threshold calibrated on ARC-C dataset. All post-softmax Top-k and Top-{greek}j employ VMC, all pre-softmax variants employ both VMC and exact SDC
Appendix F Pre- vs. post-softmax sparsification
@@
Appendix G Thresholding different layers
In this appendix, we present more results that support the decision to use denser initial layers. That is to use higher $k$ for Top-k or calibrating towards a higher $k$ in Top-{greek}j.  Figure 19 shows LLaMA2-7b model accuracy on Q&A datasets, and LLaMA2-70b model accuracy for Hellaswag dataset. The figure compares Top-k and Top-{greek}j using higher $k$ in the first 2 layers against using equal $k$ in all layers. All variants do not perform any compensations.
@@ Appendix G Thresholding different layers
Appendix H Thresholding different attention rows
@@
Appendix I Impact of GQA
In Figure 21, we show how the different attention sparsification approaches (Top-k, Top-{greek}j with and without CAPK) affect the number of required V rows. Top-k ( Figure 21(a)) guarantees exactly 128 selected elements per row of every head. However, the unified set of 4 heads in the group reaches only about 250, which indicates a certain agreement between the heads, which is mostly found in the recent tokens (as seen on the heatmap). We observed very similar characteristics in other heads and layers. Top-{greek}j approach with capping the number of selected elements to at most 128 per head yielded degraded quality of the generated text since it mainly focused attention on the most recent tokens. Finally, the ordinary Top-{greek}j in Figure 21(c), which provided good quality results, does seem to exhibit a certain variability in the number of selected elements per group, sometimes selecting more than 128 per group.
@@ Appendix I Impact of GQA
Figure 21: Attention popularity mask – LLaMA-3-8B (GQA group size$=4$), Human-eval task number 25, generative decoding iterations as rows. Left – heat map showing how many heads had the corresponding attention element in their Top-128; on the right – the number of V-rows required to be used (enough 1 head in the group is enough to require a V row).

### v2

@@ Appendix F Evaluation statistics
(a) LLaMA3-8B-Instruct HumanEval
(b) LLaMA3-70B-Instruct HumanEval
(c) LLaMA3.1-8B-Instruct LongBench-qmsum
(d) LLaMA3.1-8B-Instruct LongBench-gov_report
Figure 17: Generative Tasks - during prefill - Tradeoff between model accuracy averaged across test samples (y-axis), and the portion of required attention elements per head (x-axis). The Top-{greek}j variants employ threshold calibrated on ARC-C dataset. All post-softmax Top-k and Top-{greek}j employ VMC, all pre-softmax variants employ both VMC and exact SDC
(a) LLaMA3-8B-Instruct Human-eval
(b) LLaMA3-70B-Instruct Human-eval
(c) LLaMA3.1-8B-Instruct LongBench-qmsum
(d) LLaMA3.1-8B-Instruct LongBench-gov_report
Figure 18: Generative Tasks - during generative decoding – Tradeoff between model accuracy averaged across test samples (y-axis) and the portion of required $\bm{V}$-rows per group of heads (x-axis). The Top-{greek}j variants employ threshold calibrated on ARC-C dataset. All post-softmax Top-k and Top-{greek}j employ VMC, and all pre-softmax variants employ VMC and exact SDC. 3× to 10× reduction of $\bm{V}$ rows is achieved.
Appendix G Pre- vs. post-softmax sparsification
@@
Appendix H Thresholding different layers
In this appendix, we present more results that support the decision to use denser initial layers. That is to use higher $k$ for Top-k or calibrating towards a higher $k$ in Top-{greek}j.  Figure˜20 shows LLaMA2-7b model accuracy on Q&A datasets, and LLaMA2-70b model accuracy for Hellaswag dataset. The figure compares Top-k and Top-{greek}j using higher $k$ in the first 2 layers against using equal $k$ in all layers. All variants do not perform any compensations.
@@ Appendix H Thresholding different layers
Appendix I Thresholding different attention rows
@@
Appendix J Impact of GQA
In Figure˜22, we show how the different attention sparsification approaches (Top-k, Top-{greek}j with and without CAPK) affect the number of required $\bm{V}$ rows. Top-k ( Figure˜22(a)) guarantees exactly 128 selected elements per row of every head. However, the unified set of 4 heads in the group reaches only about 250, which indicates a certain agreement between the heads, which is mostly found in the recent tokens (as seen on the heatmap). We observed very similar characteristics in other heads and layers. Top-{greek}j approach with capping the number of selected elements to at most 128 per head yielded degraded quality of the generated text since it mainly focused attention on the most recent tokens. Finally, the ordinary Top-{greek}j in Figure˜22(c), which provided good quality results, does seem to exhibit a certain variability in the number of selected elements per group, sometimes selecting more than 128 per group.
@@
Figure 22: Attention popularity mask – LLaMA-3-8B (GQA group size=4=4), Human-eval task number 25, generative decoding iterations as rows. Left – heat map showing how many heads had the corresponding attention element in their Top-128; on the right – the number of $\bm{V}$-rows required to be used (1 head in the group is enough to require a $\bm{V}$ row).